# Probing Ranking LLMs: Mechanistic Interpretability in Information Retrieval

### v1

@@ Abstract.
Transformer networks, especially those with performance on par with GPT models, are renowned for their powerful feature extraction capabilities. However, the nature and correlation of these features with human-engineered ones remain unclear. In this study, we delve into the mechanistic workings of state-of-the-art, fine-tuning-based passage-reranking transformer networks.
Our approach involves a probing-based, layer-by-layer analysis of neurons within ranking LLMs to identify individual or groups of known human-engineered and semantic features within the network’s activations. We explore a wide range of features, including lexical, document structure, query-document interaction, advanced semantic, interaction-based, and LLM-specific features, to gain a deeper understanding of the underlying mechanisms that drive ranking decisions in LLMs.
Our results reveal a set of features that are prominently represented in LLM activations, as well as others that are notably absent. Additionally, we observe distinct behaviors of LLMs when processing low versus high relevance queries and when encountering out-of-distribution query and document sets. By examining these features within activations, we aim to enhance the interpretability and performance of LLMs in ranking tasks. Our findings provide valuable insights for the development of more effective and transparent ranking models, with significant implications for the broader information retrieval community. All scripts and code necessary to replicate our findings are made available.
@@ 1. Introduction
Motivation: For many decades, the domain of passage retrieval and reranking has predominantly used algorithms grounded in statistical human-engineered features, typically derived from the query, the document set, or their interactions. In particular, the features in training and evaluation datasets such as MSLR (Qin and Liu, 2013) are largely derived from features manually developed to support successful statistical ranking systems. However, recent advances have seen state-of-the-art passage retrieval and ranking algorithms increasingly pivot to neural network-based models. The introduction of Large Language Models (LLMs) has notably widened the performance disparity between neural-based and traditional statistical rankers.
Access to open-source ranking models, particularly those exhibiting performance comparable to GPTs, provides a unique opportunity to explore the inner mechanics of transformer-based architectures. Neural networks are recognized for their robust feature extraction capabilities. However, the representation of these features remains elusive, making it challenging to discern their nature and potential correlation with features engineered by humans. It is this difficult to develop human-understandable explanations and hard for engineers to analyze how networks are producing results.
@@
Despite their efficacy, neural networks present a challenge in terms of transparency: their feature representations are complex and the location of critical features within the network remains obscure. In this context, mechanistic interpretability (Räuker et al., 2023) aims to demystify the internal workings of transformer architectures, showing how LLMs process and learn information differently compared to traditional methods. Our research is motivated by the desire to determine whether the statistical features traditionally valued in algorithms like BM25 and tf*idf retain their relevance within LLM architectures. This study stands to bridge the gap between neural and statistical methodologies, by using probes to generate hypothesis on the inner workings of the LLMs, offering insights that could enhance the field of information retrieval by enabling more intuitive explanations and better support for model analysis and development.
@@
Summary: This study is built upon the Llama2-7b/13b architecture, specifically RankLlama (Ma et al., 2023), a LoRa fine-tuned variant of Llama-2 optimized for passage reranking tasks using the MS Marco dataset. RankLlama, a point-wise ranker, demonstrates substantial accuracy improvements over its non-LLM counterparts (Ma et al., 2023). Our analysis concentrates on extracting activations from the MLP unit of each transformer block, which is posited to contain the key feature extractors (Geva et al., 2020). We aggregate these activations for each input sequence (query-document pairs) and assign labels to these pairs corresponding to each feature that is a target of our probe. Subsequently, we employ a regression with regularization to correlate these activations with the labels. Our findings reveal a pronounced representation of certain MSLR features within the RankLlama model, while others are markedly absent. These insights allow us to formulate hypotheses concerning the underlying circuitry of ranking LLMs.
@@
Research Questions: In this study, we aim to explore several internal mechanistic aspects of ranking LLMs through probing techniques. Specifically, we seek to determine whether known statistical information retrieval (IR) features are present within the activations of LLMs. We are also interested in identifying groups of features and understanding how they may combine or interact within the LLM’s activations. Additionally, we investigate whether LLMs contain components that mimic similarity scores from models like BERT or T5. Another key question we address is whether LLMs extract the same features for highly relevant documents as they do for those with low relevance. Finally, this inquiry extends to examining whether the features extracted by LLMs remain consistent or change when the model encounters out-of-distribution queries or documents. By answering these questions, we aim to gain a deeper understanding of the inner workings of ranking LLMs and the extent to which they align with traditional IR methodologies.
@@ 1.1. Contributions and Findings
We discovered that several human-engineered metrics, such as covered query term number, covered query term ratio, min of term frequency, min and mean of stream length normalized term frequency, and variance of tf*idf, are represented in LLM feature extractors. In contrast, certain features, including sum of stream length normalized term frequency, max of stream length normalized term frequency, and BM25, show no discernible representation in LLM activations.
We identified specific combinations of MSLR features like the sum of covered query term ratio, mean strean length normalized term frequency, and variance of tf*idf), along with the sum’s square and cube, all of which exhibit a strong correlation with LLM activations.
We found that the activation patterns of the listed features remain consistent even when the model encounters out-of-distribution queries or documents on RankLlama 7b. However, they do not remain consistent for certain features of RankLlama 13b, suggesting potential overfitting during fine-tuning.
All models and datasets used in this study are open source. The scripts and code used in this study are provided along with the work.
@@ 2.1. Mechanistic Interpetability & Sparse Probing
Mechanistic Interpretability is the ambitious goal of gaining an algorithmic-level understanding of any deep neural network’s computations (Räuker et al., 2023). This goal can be achieved in numerous ways, some of them being by studying weights (e.g: weight masking (Csordás et al., 2020; Wortsman et al., 2020), continual learning (De-Arteaga et al., 2019) ), studying individual neurons (e.g: excitement based (Zhou et al., 2014), gradient-based (Ancona et al., 2019), perturbation and ablation based (Zhou et al., 2018)), by studying subnetworks (e.g: sparsity based (Meister et al., 2021), circuit analysis based (Fiacco et al., 2019)), by studying representations (e.g: tokens (Li et al., 2021), attention (Clark et al., 2019), probing (Belinkov, 2022)) etc. In this study, we focus on studying the activations of individual neurons and also some groups of neurons with the help of probing regressors. For this, we use concepts from a host of techniques mentioned above.
While the concept of reverse-engineering specific neurons within large language models (LLMs) is relatively new, existing studies such as (Geva et al., 2020, 2022) illustrate that the feed-forward layers of transformers, comprising two-thirds of the model’s parameters, function as key-value memories. These layers activate in response to specific inputs, a mechanism we aim to demystify by reverse-engineering the activation function of these neurons. A notable challenge in this endeavor is the phenomenon of superposition, where early layers in LLMs select and store a vast array of features—often exceeding the number of available neurons—in a linear combination across multiple neurons. In contrast, later layers tend to focus on more abstract features, discarding those deemed non-essential (Gurnee et al., 2023).
@@ 2.2. Inner Interpretability in IR
Most works of interpretability in information retrieval (Anand et al., 2022) has been model extrinsic (Chowdhury et al., 2023, 2024). Among model intrinsic interpretability methods, some works focus on using gradient based methods to identify important neurons (Chen et al., 2024; Fernando et al., 2019). Although gradient based methods give an accurate perspective on the flow of information, they are too fine-grained to give a human-level understanding of what is happening. A number of works have attempted to examine the inner workings of neural retrievers to understand if they satisfy IR axioms and/or to spot known features. Fan et al. (Fan et al., 2021) probe fine-tuned BERT models on three different tasks, namely document retrieval, answer retrieval and passage retrieval, to find out if these different forms of relevance lead to different modellings. Zhan et al. (Zhan et al., 2020) study the attention patterns of BERT after fine-tuning on the document ranking task and make note of how BERT dumps duplicate attention weights on high frequency tokens (like periods). Similar to them, Choi et al. (Choi et al., 2022) study attention maps in BERT and report the discovery of the IDF feature within BERT attention patterns. ColBERT investigations (Formal et al., 2021, 2022) study its term-matching mechanism and conclude that it captures a notion of term importance, which is enhanced by fine-tuning. MacAvaney et al. (MacAvaney et al., 2022) propose ABNIRML a set of diagnostic probes for neural retrieval models that allow searching for features like writing styles, factuality, sensitivity and word order in models like ColBERT and T5.
These studies, however, have been conducted in the pre-LLM era and hence are much smaller in size. It is unknown if the findings from BERT will hold over LLMs like Llama, given that BERT is an encoder-only model whereas most modern LLMs are decoder-only models. While most of the above works focus on attention heads and pattern, our work focuses on probing the MLP activation layers, which are now believed to be the primary feature extraction location within the LLM (Gurnee et al., 2023). To the best of our knowledge, this is one of the first works towards study of neuron activations in LLMs for a large scale of statistical and semantic features.
@@ 3. Identifying Context Neurons
Ranking Model: For our Ranking LLM interpretability study, we select RankLlama (Ma et al., 2023), an open source pointwise reranker that has been fine-tuned on Llama-2 with the MS MARCO dataset using LoRA. Given an input sequence (query Q, document set D), RankLlama reranks them as :
where Linear() is a linear projection layer that projects the last layer representation of the end-of sequence token to a scalar. The model is fine-tuned using a contrastive loss function. RankLlama is accessible on Huggingface111https://huggingface.co/castorini/rankllama-v1-7b-lora-passage and has demonstrated near state-of-the-art performance on the MS MARCO Dev set, as noted by Ma et al. (Ma et al., 2023). We experiment on both the 7b and 13b parameter models fine-tuned for passage reranking. At the time of writing, RankLlama seems to be the only open-source LLM fine-tuned for reranking, and thus the only choice for a mechanistic investigation. However, this study can be carried out for any fine-tuning based LLM, when they are made available in future.
Internal Architecture: We refer to the internal architecture of RankLlama in Figure 1. Each block of the Llama transformer architecture can be divided into two main groups : the Multi-Head self attention layers and the MLP layers. The Llama-2 7b and 13b architectures consist of 32 and 40 such symmetrical blocks, where each component has a dimensionality of 4096 and 5120 respectively. The feed-forward sublayer takes in the output of the multi-head attention layer and performs two linear transformations over it, with an element-wise non-linear function in between.
@@
This MLP layer applies pointwise nonlinearities to each token independently and therefore performs the majority of feature extraction for the LLM. Additionally, the MLP layers account for rd of total parameters. As a result, our study focuses on the value of the residual stream (also known as hidden state) for token at layer , right after applying the MLP layers. For further details about the transformer architecture math, we point the readers to Elhage et al.(Elhage et al., 2021).
Activation Sourcing: We set up PyTorch transformer forward hooks to mine activations from RankLlama. We tokenize and feed the query-document pairs to the pointwise LLM (as shown above) and capture activations corresponding to each input sequence across all layers. The dimension of each layer’s activation is the number of tokens times the number of hidden units in the LLM. This dimension is 4096 in Llama 7b and 5120 in Llama 13b. To reduce complexity of the computation, we aggregate activations across tokens in an input sequence within a layer. Following Gurnee et al. (Gurnee et al., 2023), we try out mean and max activation aggregation to ultimately obtain a dimensional activation vector per layer. This corresponds to aggregated activations for neurons in each layer, which is the focal point of our study. We quantize and store these activation tensors to be retrieved later for further experiments. Intuitively, we have captured the internals of the model at this point, in which we can start searching for desired features.
Target Features: The MSLR dataset (Qin and Liu, 2013) provides a comprehensive list of features that have been historically recognized as highly effective for ranking models (Han and Lei, 2018; Hochma et al., 2024). Consequently, we aim to search for a subset of these MSLR features from the RankLlama activations. Specifically, we focus on mining the following 19 features from the MSLR dataset within the LLM (“stream” means the document or passage text): Covered query term number, Covered query term ratio, Stream length, IDF (Inverse Document Frequency), Sum of term frequency, Min of term frequency, Max of term frequency, Mean of term frequency, Variance of term frequency, Sum of stream length normalized term frequency, Min of stream length normalized term frequency, Max of stream length normalized term frequency, Mean of stream length normalized term frequency, Variance of stream length normalized term frequency, Sum of tf*idf, Min of tf*idf, Max of tf*idf, Mean of tf*idf, Variance of tf*idf, BM25.
@@
Probing Datasets: To facilitate probing (Sajjad et al., 2022), we need to curate a dataset of input sequences to study the model. We select query-document pairs from the MS MARCO test set (Nguyen et al., 2016) for this purpose. For each query, we include documents that are highly relevant, highly irrelevant, and of intermediate relevance. Our input sequences for the ranking LLM consist of these query-document pairs. Note, it is important for each probing dataset to be balanced, i.e contain a uniform number of samples across the rankge of values of the feature being probed (Gurnee et al., 2023; Belinkov, 2022). For instance, when analyzing the BM25 feature, our dataset included a balanced ratio of query-documents with both low and high BM25 values. Having an unbalanced probing dataset can lead to a biased analysis as the model might have a tendency to overfit the dominant class. For each input sequence, we calculate the expected value of the feature being studied. Consequently, we compute the values for each of the 19 MSLR features 7 similarity scores for our query-document pairs.
@@
Subsequently, we perform a layer-wise analysis. For each layer and feature, we employ a sparse probe by fitting the activation vectors of the training split to the corresponding feature labels. To maintain the sparsity of the context neurons, we use Lasso (L1 regularization) with . We perform 5-fold cross validation to avoid overfitting a provide a more reliable estimate. After fitting the activations to the feature’s labels, we compute the score to measure how well the regression curve explains the variability of the labels. ranges from 0 to 1, with 1 indicating that the curve perfectly explains the varibiality in the data. A high score indicates the presence of a neuron that is sensitive to or activates for the feature being studied, whereas a negative confirms a negative correlation between the feature and the neuron’s activations.
Experimental Details: For our probing experiments, we select 500 queries from the MS MARCO dev set and retrieve the top 1000 documents for each query from the MS MARCO collections corpus using a BM25 ranker. We then compute RankLlama activations for each query-document pair using both the 7b and 13b models. The results reported in the following section are after mean aggregation and quantization for efficient storage. The size of each probing dataset is greater than 10,000 to ensure incorporation of a wide range of values for the feature being probed. The 1000 documents retrieved for a query serve as the query’s corpus for idf calculation. We use the bert-base and t5-base models for BERT and T5 score computations. We use the well-known Okapi implementation of BM25 in our experiments.

### v2

# Understanding Ranking LLMs: A Mechanistic Analysis for Information Retrieval

@@ Abstract.
Transformer networks, particularly those achieving performance comparable to GPT models, are well known for their robust feature extraction abilities. However, the nature of these extracted features and their alignment with human-engineered ones remain unexplored. In this work, we investigate the internal mechanisms of state-of-the-art, fine-tuned LLMs for passage reranking. We employ a probing-based analysis to examine neuron activations in ranking LLMs, identifying the presence of known human-engineered and semantic features. Our study spans a broad range of feature categories, including lexical signals, document structure, query-document interactions, and complex semantic representations, to uncover underlying patterns influencing ranking decisions.
Through experiments on four different ranking LLMs, we identify statistical IR features that are prominently encoded in LLM activations, as well as others that are notably missing. Furthermore, we analyze how these models respond to out-of-distribution queries and documents, revealing distinct generalization behaviors. By dissecting the latent representations within LLM activations, we aim to improve both the interpretability and effectiveness of ranking models. Our findings offer crucial insights for developing more transparent and reliable retrieval systems, and we release all necessary scripts and code to support further exploration.
@@ 1. Introduction
For many decades, the domain of passage retrieval and reranking has predominantly used algorithms grounded in statistical human-engineered features, typically derived from the query, the document set, or their interactions. In particular, the features in training and evaluation datasets such as the learning-to-rank MSLR dataset (Qin and Liu, 2013) are largely derived from features manually developed to support successful statistical ranking systems. However, recent advances have seen state-of-the-art passage retrieval and ranking algorithms increasingly pivot to neural network-based models. The introduction of Large Language Models (LLMs) has notably widened the performance disparity between neural-based and traditional statistical rankers.
@@
Access to open-source ranking models, particularly those exhibiting performance comparable to GPTs, provides a unique opportunity to explore the inner workings of transformer-based architectures. Neural networks are recognized for their robust feature extraction capabilities; however, the representation of these features remains elusive, making it challenging to discern their nature and potential correlation with features engineered by humans. As a result, despite their efficacy, neural networks present a challenge in terms of transparency: their feature representations are complex and the location of critical features within the network remains obscure. In this context, mechanistic interpretability (Räuker et al., 2023) aims to demystify the internal workings of transformer architectures, showing how LLMs process and learn information differently compared to traditional methods. Our research is motivated by the desire to determine whether the statistical features traditionally valued in algorithms like BM25 and tf*idf are somehow encoded within LLM architectures. This study stands to bridge the gap between neural and statistical methodologies, by using probes to generate hypotheses on the inner workings of the LLMs, offering insights that could enhance the field of information retrieval by enabling more intuitive explanations and better support for model design and analysis.
@@
Summary: This study is built upon four different LLM architectures : Llama2-7b, Llama2-13b, Llama3.1-8b, and Pythia-6.9b. For each of these we use or create a LoRa fine-tuned variant (e.g., RankLlama (Ma et al., 2023)), optimized for passage reranking tasks using the MS Marco dataset. These point-wise rankers demonstrate substantial accuracy improvements over their non-LLM counterparts. Our analysis concentrates on extracting activations from the MLP unit of each transformer block, which is posited to contain the key feature extractors (Geva et al., 2020). We aggregate these activations for each input sequence (query-document pairs) and assign labels to these pairs corresponding to each feature that is a target of our probe. Subsequently, we employ a regression with regularization to correlate these activations with the labels. Our findings reveal a pronounced representation of certain MSLR features within the ranking LLMs, while others are markedly absent. We also observe that distinct LLM architectures capture very similar statistical IR features, when fine-tuned using the same dataset. These insights will allow future researchers to formulate hypotheses concerning the underlying circuitry of ranking LLMs.
@@
Research Questions: In this study, we aim to explore several internal mechanistic aspects of ranking LLMs through probing techniques. Specifically, we seek to determine whether known statistical information retrieval (IR) features are present within the activations of LLMs. We are also interested in identifying groups of features and understanding how they may combine or interact within the LLM’s activations. Additionally, we investigate whether LLMs contain components that mimic similarity scores from models like BERT or T5. It is also of interest to find out if different LLM architectures encode the same statistical IR features within their activations. Finally, this inquiry extends to examining whether the latent features encoded within activations of LLMs remain consistent or change when the model encounters out-of-distribution queries or documents. By answering these questions, we aim to gain a deeper understanding of the inner workings of ranking LLMs and the extent to which they align with traditional IR methodologies.
@@ 1.1. Contributions and Findings
We discovered that several human-engineered metrics, such as covered query term number, covered query term ratio, min of term frequency, min and mean of stream length normalized term frequency, and variance of tf*idf, are prominently encoded within LLM activations. In contrast, certain features, including sum of stream length normalized term frequency, max of stream length normalized term frequency, and BM25, show no discernible representation in LLM activations.
We fine-tuned and probed two additional reranking models, Llama3.1-8b and Pythia-6.9b, resulting in RankLlama3-8b and RankPythia, alongside probing RankLlama2-7b and RankLlama2-13b. The results revealed broadly consistent outcomes across all models, suggesting that different LLM architectures, when fine-tuned similarly, encode latent features in a comparable way within their activations.
We found that the activation patterns of the listed features remain consistent even when the model encounters out-of-distribution queries or documents on RankLlama2-7b, RankLlama3-8b and RankPythia-6.9b. However, they do not remain consistent for certain features of RankLlama2-13b, suggesting potential overfitting in the LLM during fine-tuning.
We identified specific combinations of MSLR features, like the sum of covered query term ratio, mean stream length normalized term frequency, and variance of tf*idf, along with the sum’s square and cube, all of which also exhibit a strong correlation with LLM activations.
Finally, we show further evidence on the presence of these encoded features using Shapley values. All scripts, datasets, and fine-tuned ranking LLMs developed and/or used in this study are made publicly available.111https://suppressed.for.review
@@ 2.1. Mechanistic Interpetability & Sparse Probing
Mechanistic Interpretability is the ambitious goal of gaining an algorithmic-level understanding of any deep neural network’s computations (Räuker et al., 2023). This goal can be achieved in numerous ways, some of them being by studying weights (e.g., weight masking (Csordás et al., 2020; Wortsman et al., 2020), by continual learning (De-Arteaga et al., 2019)), by studying individual neurons (e.g., excitement based (Zhou et al., 2014), activation patching (Makelov et al., 2023), gradient-based (Ancona et al., 2019), perturbation and ablation based (Zhou et al., 2018)), by studying subnetworks (e.g., sparsity based (Meister et al., 2021), circuit analysis based (Fiacco et al., 2019)), by studying representations (e.g., tokens (Li et al., 2021), attention (Clark et al., 2019), probing (Belinkov, 2022; Gurnee and Tegmark, 2023; Chen et al., 2023)), by training sparse auto encoders (Makelov et al., 2024; Kissane et al., 2024; Cunningham et al., 2023) etc. In this study, we focus on studying the activations of individual neurons and also some groups of neurons with the help of probing regressors. To that end, we use concepts from a host of techniques mentioned above.
While the concept of reverse-engineering specific neurons within large language models (LLMs) is relatively new, existing studies (Geva et al., 2020, 2022) illustrate that the feed-forward layers of transformers, comprising two-thirds of the model’s parameters, function as key-value memories. These layers activate in response to specific inputs, a mechanism we aim to demystify by reverse-engineering the activation function of these neurons. A notable challenge in this endeavor is the phenomenon of superposition, where early layers in LLMs select and store a vast array of features—often exceeding the number of available neurons—in a linear combination across multiple neurons. In contrast, later layers tend to focus on more abstract features, discarding those deemed non-essential (Gurnee et al., 2023).
@@
| Model Name | Base Model | DEV MRR@10 | DL19 NDCG@10 | DL20 NDCG@10 |
| BM25 | Lucene (Katsimpras and Paliouras, 2024) | 16.7 | 50.6 | 48.0 |
[…]
| RankLLama2-13b | Llama2 (Touvron et al., 2023) | 45.2 | 76.0 | 77.9 |
| RankLlama3-8b | Llama3 (Dubey et al., 2024) | 42.8 | 74.3 | 72.4 |
| RankPythia-6.9b | Pythia (Biderman et al., 2023) | 42.9 | 76.0 | 71.5 |
@@ 2.2. Inner Interpretability in IR
Most works of interpretability in information retrieval (Anand et al., 2022) have been model extrinsic (Rahimi et al., 2021; Singh and Anand, 2019; Chowdhury et al., 2023). Among model intrinsic interpretability methods, some works focus on using gradient-based methods to identify important neurons (Chen et al., 2024; Fernando et al., 2019). Although gradient based methods give an accurate perspective on the flow of information, they are too fine-grained to give a human-level understanding of the LLM’s inner circuit (Anand et al., 2022). A number of works have attempted to examine the inner workings of neural retrievers to understand if they satisfy IR axioms and/or to spot known features. Fan et al. (Fan et al., 2021) probe fine-tuned BERT models on three different tasks, namely document retrieval, answer retrieval, and passage retrieval, to find out if these different forms of relevance lead to different models. Zhan et al. (Zhan et al., 2020) study the attention patterns of BERT after fine-tuning on the document ranking task and make note of how BERT dumps duplicate attention weights on high frequency tokens (like periods). Similar to them, Choi et al. (Choi et al., 2022) study attention maps in BERT and report the discovery of the IDF feature within BERT attention patterns. ColBERT investigations (Formal et al., 2021, 2022) study its term-matching mechanism and conclude that it captures a notion of term importance, which is enhanced by fine-tuning. MacAvaney et al. (MacAvaney et al., 2022) propose ABNIRML a set of diagnostic probes for neural retrieval models that allow searching for features like writing styles, factuality, sensitivity and word order in models like ColBERT and T5. Recently Parry et al. (Parry et al., 2024) provide a framework for diagnostic analysis and intervention within Neural Ranking models by generating perturbed inputs followed by activation patching.
These studies, however, have been conducted in the pre-LLM era and hence are much smaller in size. It is unknown if the findings from BERT will carry over to LLMs like Llama, given that BERT is an encoder-only model whereas most modern LLMs are decoder-only models. While most of the above works focus on attention heads and pattern, our work focuses on probing the MLP activation layers, which are now believed to be the primary feature extraction location within the LLM (Gurnee et al., 2023). To the best of our knowledge, this is one of the first works towards study of neuron activations in LLMs for a large scale of statistical and semantic features.
@@ 3. Identifying Context Neurons
Ranking Models: For our Ranking LLM interpretability study, we first select RankLlama (Ma et al., 2023), an open source pointwise reranker that has been fine-tuned on Llama-2 with the MS MARCO dataset using LoRA. Given an input sequence (query Q, document set D), RankLlama reranks them as :
where Linear() is a linear projection layer that projects the last layer representation of the end-of sequence token to a scalar. The model is fine-tuned using a contrastive loss function. RankLlama is accessible on Huggingface222https://huggingface.co/castorini/rankllama-v1-7b-lora-passage and has demonstrated near state-of-the-art performance on the MS MARCO Dev set, as noted by Ma et al. (Ma et al., 2023). We experiment on both the 7b and 13b parameter models fine-tuned for passage reranking.
To facilitate comparison with other ranking architectures, we additionally fine-tune Llama3.1-8b and Pythia-6.9b using an approach similar to the fine-tuning methodology employed by the RankLlama authors. Each of these LLMs is fine-tuned with a LoRa rank of 32 for 1 epoch, and their performance is evaluated on the commonly used TREC DL19 and DL20 datasets. Table 1 compares the ranking performance of all LLMs used in this study, indicating that, once fine-tuned, they achieve similar effectiveness.
Internal Architecture: We refer to the internal architecture of RankLlama2-7b in Figure 1. Each block of the Llama transformer architecture can be divided into two main groups: the Multi-Head self attention layers and the MLP layers. For example, the Llama-2 7b and 13b architectures consist of 32 and 40 such identical blocks, where each component has a dimensionality of 4096 and 5120 respectively. The feed-forward sublayer takes in the output of the multi-head attention layer and performs two linear transformations over it, with an element-wise non-linear function in between.
@@
This MLP layer applies pointwise nonlinearities to each token independently and therefore performs the majority of feature extraction for the LLM. Additionally, the MLP layers account for rd of total parameters. As a result, our study focuses on the value of the residual stream (also known as hidden state) for token at layer , right after applying the MLP layers. Elhage et al. (Elhage et al., 2021) provide further details about the transformer architecture math.
Activation Sourcing: We set up PyTorch transformer forward hooks to mine activations from the ranking LLMs. We tokenize and feed the query-document pairs to the pointwise LLM discussed above, and capture activations corresponding to each input sequence across all layers. The dimension of each layer’s activation is the number of tokens times the number of hidden units in the LLM. This dimension is 4096 in Llama2-7b and 5120 in Llama2-13b. To reduce complexity of the computation, we aggregate activations across tokens in an input sequence within a layer. Following Gurnee et al. (Gurnee et al., 2023), we try out mean and max activation aggregation to ultimately obtain a dimensional activation vector per layer. This corresponds to aggregated activations for neurons in each layer, which is the focal point of our study. We quantize and store these activation tensors to be used later for further experiments. Intuitively, we have captured the internals of the model at this point, in which we can start searching for desired features.
Target Features: The MSLR dataset (Qin and Liu, 2013) provides a comprehensive list of features that have been historically recognized as highly effective for ranking models (Han and Lei, 2018; Hochma et al., 2024). Consequently, we aim to search for a subset of these MSLR features within the LLM activations. Specifically, we focus on mining the following 19 features from the MSLR dataset within the LLM (“stream” means the document or passage text): Min TF, Min(TF/L), Min TF*IDF, Covered QT Number, Covered QT Ratio, Mean(TF/L), Variance TF*IDF, BM25, Mean TF*IDF, Variance(TF/L), Mean TF, Variance TF, Sum TF, Max TF, Max(TF/L), Stream length, Sum TF, Max TF*IDF and Sum TF*IDF.
@@
Probing Datasets: To facilitate probing (Sajjad et al., 2022), we need to curate a dataset of input sequences to study the model. We select query-document pairs from the MS MARCO test set (Nguyen et al., 2016) for this purpose. For each query, we include documents that are highly relevant, highly irrelevant, and of intermediate relevance. Our input sequences for the ranking LLM consist of these query-document pairs. Note that it is important for each probing dataset to be balanced – i.e., contain a uniform number of samples across the range of values of the feature being probed (Gurnee et al., 2023; Belinkov, 2022). For instance, when analyzing the BM25 feature, our dataset included a balanced ratio of query-documents with both low and high BM25 values. Having an unbalanced probing dataset can lead to a biased analysis as the model might have a tendency to overfit the dominant class. For each input sequence, we calculate the expected value of the feature being studied. Consequently, we compute the values for each of the 19 MSLR features and 7 similarity scores for our query-document pairs.
@@
Subsequently, we perform a layer-wise analysis. For each layer and feature, we employ a sparse probe by fitting the activation vectors of the training split to the corresponding feature labels. To maintain the sparsity of the context neurons, we use Lasso (L1 regularization) with . We perform 5-fold cross validation to avoid overfitting and provide a more reliable estimate. After fitting the activations to the feature’s labels, we compute the score to measure how well the regression curve explains the variability of the labels. ranges from 0 to 1, with 1 indicating that the curve perfectly explains the varibiality in the data. A high score indicates the presence of a neuron that is sensitive to or activates for the feature being studied, whereas a negative confirms a negative correlation between the feature and the neuron’s activations.
Experimental Details: For our probing experiments, we select 500 queries from the MS MARCO dev set and retrieve the top 100 documents for each query from the MS MARCO collections corpus using a BM25 ranker. We then compute activations corresponding to each query-document pair using all four models. The results reported in the following section are after mean aggregation and quantization for efficient storage. The size of each probing dataset is greater than 5,000 to ensure incorporation of a wide range of values for the feature being probed. The 100 documents retrieved for a query serve as the query’s corpus for idf calculation. We use the bert-base and t5-base models for BERT and T5 score computations. We use the well-known Okapi implementation of BM25 in our experiments.